OpenStudy (anonymous):

When using the substitution method, what expression for x in the first equation would be substituted in the second equation? x + 3y = –7 2x + 4y = –4 3y – 7 –3y – 7 3y + 7 –3y + 7

OpenStudy (anonymous):

@85295james

OpenStudy (anonymous):

x=8 y=-5

OpenStudy (anonymous):

would it be the second choice

OpenStudy (anonymous):

i think the first one ??????????????? sorry

hartnn (hartnn):

from first equation, x+3y=-7 subtract 3y from both sides, so x+3y-3y=-3y-7 x+0=-3y-7 x=-3y-7 so 2nd choice :)

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

could u help me with this one Sam and Chris went to “Lots O Fun” to play laser tag and video games for Chris’s birthday. Sam played 3 games of laser tag, 5 video games, and spent $17 total. Chris played 4 games of laser tag, 7 video games, and spent $23 total. How much does one video game cost to play? $1 $2 $3 $4

hartnn (hartnn):

let laser tag cost x $ and video game cost y $. so u can form the equations : 3x+5y = 17 4x+7y = 23 just solve these 2 equations and find the value of y. that would be your answer.

OpenStudy (anonymous):

$2

OpenStudy (anonymous):

oh ok'

hartnn (hartnn):

lol its $ 1 when u solve :P

OpenStudy (anonymous):

o $hit sorry

OpenStudy (anonymous):

i got the same answer

OpenStudy (anonymous):

hehehe

OpenStudy (anonymous):

i have one more 2x – 7y = 1 x – 8y = –4 (2, 4) (4, 2) (1, 4) (4, 1)

OpenStudy (anonymous):

4,1

hartnn (hartnn):

yup :)

OpenStudy (anonymous):

can you just open a new one i love medals lol

OpenStudy (anonymous):

yaya

OpenStudy (anonymous):

cool